OpenStudy (anonymous):

how can you divide 1 into 9 if you havn't found out yet do this (9 X 8)31 X 12=????

OpenStudy (anonymous):



OpenStudy (anonymous):

@baking4lyf

OpenStudy (anonymous):

is the 31 an exponent or something

OpenStudy (midhun.madhu1987):

can u draw the question or post the screenshot of the question?

OpenStudy (anonymous):

(9 X 8)^31 X 12=??? follow PEMDAS, the order of operations. look it up if u need help with it. (72)^31= 3.77 3.77 * 12= 45.24

OpenStudy (anonymous):

do you have answer choices

OpenStudy (anonymous):

no just explain

OpenStudy (anonymous):

oh, i thought that above was ur problem^^

OpenStudy (anonymous):

-_- well bye now O_O bye